OpenStudy (shamil98):

If the value of \[\large 34 \times \frac{ \int\limits_{0}^{\pi}(\sin x)^{3\sqrt{2}+3} dx }{ \int\limits_{0}^{\frac{ \pi }{ 2 }} (\cos x)^{3\sqrt{2} -3} dx } = a \sqrt{2} + b\] for integers a,b. \[\ \text{What is the value of} ~~~a + b ~~~?\]

OpenStudy (usukidoll):

the value is I quit...this is crazy

OpenStudy (shamil98):

lol xD

OpenStudy (anonymous):

Might want to use McLaurin series.

OpenStudy (anonymous):

How accurate of an answer do you want? You could use a series expansion or other approximate method (Brute force numerical methods, etc.)

OpenStudy (anonymous):

Grabs pencil + piece of paper, cracks knuckles...here we go..

OpenStudy (anonymous):

\[ \sin(x) = \sum_{k=0}^{\infty}\frac{(-1)^k}{(2k+1)}x^{2k+1} \]

OpenStudy (anonymous):

I'm not sure how much this will help exactly.

OpenStudy (nincompoop):

wio is giving us more work! laughing out loud jaykay

OpenStudy (anonymous):

Besides the constraint on a and b being integers, what other information is available to solve for two unknowns with only one equation?

OpenStudy (anonymous):